What is the pH of rainwater in pristine environments? Assume the temperature is 25°C, the total pressure is 1 atm and the rainwater is in equilibrium with the atmosphere.

Solution: 5 species are present in the aqueous phase H2CO3(aq), HCO3-(aq), CO3-2(aq), H+ and OH-

Equations: 1H2CO3 → H+ + HCO3-

(carbonic acid dissociation)

K1 = [H+][HCO3-]/[H2CO3] = 4.47 x 10-7M
Answer #1

According to Henry's law: [CO2]in water = KH * PCO2 = 3.4*10-2 mol.atm/L * 1 atm = 3.4*10-2 M

Ka1 = 4.45*10-7 = [H+][HCO3- ]/[CO2] = [H+]2/(3.4*10-2)

i.e. [H+] = 1.23*10-4

i.e. pH = -Log[H+] = -Log(1.23*10-4)

Therefore, the pH of rainwater under given conditions = 3.91